The Linux HOWTO Index

by Tim Bynum, linux-howto@metalab.unc.edu v2.11.15, 26 October 1999

This document contains an index to the Linux HOWTOs and mini-HOWTOs, as well as other information about the HOWTO project.

Contents

1 What Are Linux HOWTOs? 1

2 Where Can I Get Linux HOWTOs? 1 2.1 HOWTO Translations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Index 2 3.1 HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.2 mini-HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.3 Special HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.4 Unmaintained HOWTOs and mini-HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

4 Writing and Submitting a HOWTO 15

5 Copyright 16

1 What Are Linux HOWTOs?

Linux HOWTOs are documents which describe in detail a certain aspect of configuring or using Linux. For example, there is the Installation HOWTO, which gives instructions on installing Linux, and the Mail HOWTO, which describes how to set up and configure mail under Linux. Other examples include the NET-3 HOWTO and the Printing HOWTO. HOWTOs are comprehensive docs - much like an FAQ but generally not in question-and-answer format. However, many HOWTOs contain an FAQ section at the end. There are several HOWTO formats available: plain text, PostScript, DVI, and HTML. In addition to the HOWTOs, there are a multitude of mini-HOWTOs on short, specific subjects. They are only available in plain text and HTML format.

2 Where Can I Get Linux HOWTOs?

HOWTOs can be retrieved via anonymous FTP from the following sites:

   • ftp://metalab.unc.edu/pub/Linux/docs/HOWTO
   • ftp://tsx-11.mit.edu/pub/linux/docs/HOWTO

as well as many mirror sites. You can also browse HOWTOs in HTML format <http://metalab.unc.edu/LDP/HOWTO/>. Many mirror

sites <http://metalab.unc.edu/LDP/hmirrors.html> mirror the HTML files.

metalab.unc.edu is heavily used, so please use a mirror site if possible. HOWTOs are also posted towards the beginning of the month to the Usenet newsgroup

comp.os.linux.answers. There is a tool called NewstoHOWTO <ftp://metalab.unc.edu/pub/Linux/

system/news/misc/> that will assemble the postings.

2.1 HOWTO Translations

HOWTO translations are available on metalab.unc.edu <ftp://metalab.unc.edu/pub/Linux/docs/HOWTO/ translations> and mirrors around the world. So far there are:

• Chinese (zh)

• Croatian (hr)

• French (fr)

• German (de)

• Hellenic (el)

• Indonesian (id)

• Italian (it)

• Japanese (ja)

• Korean (ko)

• Polish (pl)

• Slovenian (sl)

• Spanish (es)

• Swedish (sv)

• Turkish (tr) If you know of any other translation projects, please let me know and I will add them to this list. If you are interested in getting your translations archived on metalab.unc.edu, please read the directory structure specification at http://metalab.unc.edu/pub/Linux/docs/HOWTO/translations/Directory-Structure and get in touch with me.

3 Index

3.1 HOWTOs

The following Linux HOWTOs are currently available:

• 3Dfx HOWTO, by Bernd Kreimeier <bk@gamers.org>. How to use 3Dfx graphics accelerator chip support. Updated 6 February 1998.

• AI-Alife HOWTO, by John A. Eikenberry <jae@ai.uga.edu>. Information about AI software for Linux. Updated 20 October 1999.

• AX25 HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. How to configure AX25 networking for Linux. Updated 17 October 1997.

• Access HOWTO, by Michael De La Rue <access-howto@ed.ac.uk>. How to use adaptive technology with Linux. Updated 28 March 1997.

• Alpha HOWTO, by David Mosberger <davidm@azstarnet.com>. Overview of Alpha systems and processors. Updated 6 June 1997.

• Assembly HOWTO, by Konstantin Boldyshev <fare@tunes.org> François-René Rideau tunes.org>. Information on programming in x86 assembly. Updated 19 Sep 1999.

• Bash Prompt HOWTO, by Giles Orr <giles@interlog.com>. Creating and controlling terminal and xterm prompts. Updated 07 January 1999.

• Belgian HOWTO, by Dag Wieers <dag@mind.be>. Information on using Linux with Belgian-specific features. Updated 07 January 1999.

• Benchmarking HOWTO, by André D. Balsa <andrewbalsa@usa.net>. How to do basic benchmarking. Updated 15 August 1997.

• Beowulf HOWTO, by Jacek Radajewski <jacek@usq.edu.au> Douglas Eadline plogic.com>. Introduces the Beowulf Supercomputer architecture and provides background information on parallel programming. Updated 22 November 1998.

• BootPrompt HOWTO, by Paul Gortmaker <gpg109@rsphy1.anu.edu.au>. List of boot time arguments and overview of booting software. Updated May 15, 1999.

• Bootdisk HOWTO, by Tom Fawcett <fawcett@croftj.net>. How to create a boot/root maintenance disk for Linux. Updated May 15, 1999.

• Busmouse HOWTO, by Chris Bagwell <cbagwell@sprynet.com>. Information on bus mouse compatibility with Linux. Updated 15 June 1998.

• CD Writing HOWTO, by Winfried Trümper <winni@xpilot.org>. How to write CDs. Updated 19 September 1999.

• CDROM HOWTO, by Jeff Tranter <jeff_tranter@pobox.com>. Information on CDROM drive compatibility for Linux. Updated 24 March 1999.

• CVS RCS HOWTO, by Alavoor Vasudevan <alavoor@yahoo.com>. This document is a practical guide to very quickly setup CVS/RCS. Updated 23 October 1999.

• Chinese HOWTO, by Chih-Wei Huang <cwhuang@phys.ntu.edu.tw>. How to configure Linux for use with the Chinese characterset. Updated 2 June 1998.

• Commercial HOWTO, by Mr. Poet <poet@linuxports.com>. Listing of commercial software products for Linux. Updated 7 March 1999.

• Config HOWTO, by Guido Gonzato <guido@ibogfs.cineca.it>. How to fine-tune and customize your Linux system. Updated 6 September 1999.

• Consultants HOWTO, by Mr. Poet <poet@linuxports.com>. Listing of Linux consultants. Updated 12 September 1999.

• Cyrillic HOWTO, by Alexander L. Belikoff <abel@bfr.co.il>. How to configure Linux for use with the Cyrillic characterset. Updated 23 January 1998.

• DNS HOWTO, by Nicolai Langfeldt <janl@math.uio.no>. How to set up DNS. Updated 11 February 1999.

• DOS/Win to Linux HOWTO, by Guido Gonzato <guido@ibogfs.cineca.it>. How to move from DOS/Windows to Linux. Updated 22 February 1999.

• DOSEMU HOWTO, by Uwe Bonnes <bon@elektron.ikp.physik.th-darmstadt.de> David Hodges erewhon.ak.planet.co.nz>. HOWTO about the Linux MS-DOS Emulator, DOSEMU. Updated dosemu-0.98.8, 24 April 1999.

• Danish HOWTO, by Niels Kristian Bech Jensen <nkbj@image.dk>. How to configure Linux for use with the Danish characterset. Updated 20 February 1999.

• Diskless HOWTO, by Alavoor Vasudevan <alavoor@yahoo.com>. How to set up a diskless Linux box. Updated 21 October 1999.

• Distribution HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. A list of Linux distributions. Updated 8 May 1999.

• Ecology HOWTO, by Werner Heuser <wehe@snafu.de>. Discusses ways Linux computers can be used as a mean to protect our environment, by using its features to save power or paper. Updated 17 September 1999 .

• Emacs Beginner’s HOWTO, by Jeremy D. Zawodny <jzawodn@wcnet.org>. This document introduces Linux users to the Emacs editor. Updated .

• Emacspeak HOWTO, by Jim Van Zandt <jrv@vanzandt.mv.com>. How to use ’emacspeak’ with Linux. Updated 10 April 1999.

• Esperanto HOWTO, by Wolfram Diestel <diestel@rzaix340.rz.uni-leipzig.de>. How to use Esperanto in general and ISO-8859-3 in special with Linux. Updated August 1999.

• Ethernet HOWTO, by Paul Gortmaker <gpg109@rsphy1.anu.edu.au>. Information on Ethernet hardware compatibility for Linux. Updated 5 May 1999.

• Finnish HOWTO, by Pekka Taipale <pjt@iki.fi>. How to configure Linux for use with the Finnish characterset. Updated 14 February 1996.

• Firewall HOWTO, by Mark Grennan <markg@netplus.net>. How to set up a firewall using Linux. Updated 26 September 1999.

• Framebuffer HOWTO, by Alex Buell <alex.buell@tahallah.demon.co.uk>. This document describes how to use the framebuffer devices in Linux. Updated 22 July 1999.

• Ftape HOWTO, by Kevin Johnson <kjj@pobox.com>. Information on ftape drive compatibility with Linux. Updated August 1998.

• GCC HOWTO, by Daniel Barlow <daniel.barlow@linux.org>. How to set up the GNU C compiler and development libraries. Updated 28 February 1996.

• German HOWTO, by Winfried Trümper <winni@xpilot.org>. Information on using Linux with German-specific features. Updated 19 March 1997.

• Glibc2 HOWTO, by Eric Green <ejg3@cornell.edu>. How to install and migrate to the glibc2 library. Updated 22 June 1998.

• HAM HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. HOWTO configure amateur radio software for Linux. Updated 1 April 1997.

• HOWTO HOWTO, by Mark F. Komarinski <markk@cgipc.com>. Getting a new LDP author up and running with tools, ideas, and. Updated 1 April 1997.

• HOWTO Index, by Tim Bynum <linux-howto@metalab.unc.edu>. Index of HOWTO documents about Linux. Updated 26 October 1999.

• Hardware Compatibility HOWTO, by Patrick Reijnen <antispam.patrickr@antispam.bart.nl>. A list of hardware known to work with Linux. Updated 28 September 1999.

• Hebrew HOWTO, by Yair G. Rajwan <yair@hobbes.jct.ac.il>. How to configure Linux for use with the Hebrew characterset. Updated 12 September 1995.

• INFO-SHEET, by Michael K. Johnson <johnsonm@redhat.com>. Generic introduction to the Linux operating system. Updated 1 September 1998.

• IP Masquerade HOWTO, by Ambrose Au <ambrose@writeme.com>. This document describes how to enable the Linux IP Masquerade feature on a given Linux host. Updated 21 October 1999.

• IPCHAINS HOWTO, by Paul Russell <Paul.Russell@rustcorp.com.au>. Install and configure the enhanced IP firewalling chains software. Updated 12 March 1999.

• IPX HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. How to install and configure IPX networking. Updated 06 May 1998.

• IR HOWTO, by Werner Heuser <r2d2c3po@zedat.fu-berlin.de>. An introduction to the software provided by the Linux/IR project. Updated 20 September 1999.

• ISP Hookup HOWTO, by Egil Kvaleberg <egil@kvaleberg.no>. Basic introduction to hooking up to an ISP. Updated 5 March 1998.

• Installation HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. How to obtain and install Linux. Updated 26 September 1999.

• Intranet Server HOWTO, by Pramod Karnad <karnad@indiamail.com>. How to setup a Linux Intranet server. Updated 7 August 1997.

• Italian HOWTO, by Marco “Gaio” Gaiarin <gaio@dei.unipd.it>. How to configure Linux for use with the Italian characterset. Updated 3 November 1998.

• Java-CGI HOWTO, by David H. Silber <dhs@orbits.com>. How to set up Java-capable CGI bin. Updated 1 December 1998.

• Kernel HOWTO, by Brian Ward <ward@blah.math.tu-graz.ac.at>. Upgrading and compiling the Linux kernel. Updated 5 June 1999.

• Keyboard and Console HOWTO, by Andries Brouwer <aeb@cwi.nl>. Information about the Linux keyboard, console and non-ASCII characters. Updated 25 February 1998.

• KickStart HOWTO, by Martin Hamilton <martinh@gnu.org>. Briefly describes how to use the RedHat Linux KickStart system to rapidly install large numbers of identical Linux boxes. Updated 11 January 1999.

• Kiosk HOWTO, by Gene Wilburn <genew@rom.on.ca>. This document provides a guide for setting up a WWW-based kiosk using Linux, X11R6, FVWM2, Netscape Navigator 4.X, and a customized trackball. Updated October 1999.

• LDAP HOWTO, by Luiz Ernesto Pinheiro Malere <malere@yahoo.com>. Information about installing, configuring, running and maintaining an LDAP Server on Linux. Updated 20 June 1999.

• Laptop HOWTO, by Werner Heuser <wehe@snafu.de>. Contains a concise survey of laptop related documents laptop related Linux features. Updated 2 October 1999.

• LinuxDoc+Emacs+Ispell HOWTO, by Philippe Martin <feloy@wanadoo.fr>. Assist writers and translators of Linux HOWTOs or any other paper for the Linux Documentation Project. Updated 27 February 1998.

• Loopback Encrypted Filesystem HOWTO, by Ryan T. Rhea <zzrhear@pobox.winthrop.edu>. This document explains how to setup and then use a filesystem that, when mounted by a user, dynamically and transparently encrypts its contents. Updated 22 September 1999.

• META-FAQ, by Michael K. Johnson <johnsonm@redhat.com>. A listing of Linux sources of information. Updated 25 October 1997.

• MGR HOWTO, by Vincent Broman <broman@nosc.mil>. Information on the MGR graphics interface for Linux. Updated 30 May 1996.

• MILO HOWTO, by David A. Rusling <david.rusling@reo.mts.dec.com>. How to use the Alpha Linux Miniloader (MILO). Updated 6 December 1996.

• MIPS HOWTO, by Ralf Baechle <ralf@gnu.org>. Describes the MIPS port of the Linux operating system, common problems and their solutions, availability and more. Updated 31 March 1999.

• MP3 HOWTO, by Philip Kerr <phil@websentric.com>. This document describes the hardware, software, and procedures needed to encode and play MP3 sound files under Linux. Updated September 1999.

• Mail HOWTO, by Guylhem Aznar <guylhem@danmark.linux.eu.org>. Information on electronic mail servers and clients. Updated July 1999.

• Mail User HOWTO, by Eric S. Raymond <esr@thyrsus.com>. This document is an introduction to the world of electronic mail. Updated 07 May 1999.

• Modem HOWTO, by David S. Lawyer <bf347@lafn.org>. Help with selecting, connecting, configuring, trouble-shooting, and understanding modems for a PC. Updated October 1999.

• Multi-Disk HOWTO, by Stein Gjoen <sgjoen@nyx.net>. How to setup multiple hard disk drives. Updated 27 May 1999.

• Multicast HOWTO, by Juan-Mariano de Goyeneche <jmseyas@dit.upm.es>. This HOWTO tries to cover most aspects related to multicast over TCP/IP networks. Updated 20 March 1998.

• Mutt GnuPG PGP HOWTO, by Andrés Seco <AndresSH@ctv.es>, J.Horacio M.G. ciberia.es>. This document briefly explains how to configure Mutt-i, PGP and GnuPG. Updated 2 July 1999.

• NET3-4 HOWTO, by Mr. Poet <poet@linuxports.com>. Howto configure TCP/IP networking under Linux. Updated August 1999.

• NFS HOWTO, by Nicolai Langfeldt <janl@math.uio.no>. How to set up NFS clients and servers. Updated 1 October 1999.

• NIS HOWTO, by Thorsten Kukuk <kukuk@suse.de>. Information on using NIS/YP on Linux systems. Updated 9 March 1999.

• Networking Overview HOWTO, by Daniel López Ridruejo <ridruejo@esi.us.es>. The purpose of this document is to give an overview of the networking capabilities of the Linux Operating System providing pointers for further information and implementation details. Updated 17 September 1999.

• Online Troubleshooting Resources HOWTO, by Howard Mann <howardm@xmission.com>. This document will direct Linux users to resources available on the Internet. Updated 18 September 1999.

• Optical Disk HOWTO, by Skip Rye <Skip_Rye@faneuil.com>. How to use optical disk drives with Linux. Updated 11 December 1998.

• Oracle HOWTO, by Paul Haigh <paul@nailed.demon.co.uk>. How to setup Oracle as a database server. Updated 4 August 1998.

• PCI HOWTO, by Michael Will <Michael.Will@student.uni-tuebingen.de>. Information on PCIarchitecture compatibility with Linux. Updated 30 March 1997.

• PCMCIA HOWTO, by Dave Hinds <dhinds@allegro.stanford.edu>. How to install and use PCMCIA Card Services. Updated 15 May 1999.

• PLIP Install HOWTO, by Gilles Lamiral <lamiral@mail.dotcom.fr>. This paper describes how to install a Gnu|Linux distribution on a computer without an Ethernet card, nor cdrom, but just a local floppy drive and a nfs server attached by a Null-Modem parallel cable. Updated 15 May 1999.

• PPP HOWTO, by Robert Hart <hartr@interweft.com.au>. Information on using PPP networking with Linux. Updated 31 March 1997.

• PalmOS HOWTO, by David H. Silber <pilot@orbits.com>. How to use your Palm OS device with a Linux system. Updated 20 September 1998.

• Parallel Processing HOWTO, by Hank Dietz <pplinux@ecn.purdue.edu>. Discussion of parallel processing approaches for Linux. Updated 5 January 1998.

• Plug and Play HOWTO, by David Lawyer <bf347@lafn.org>. How to get your Linux system to support Plug-and-Play. Updated October 1999.

• Polish HOWTO, by Sergiusz Pawlowicz <ser@arch.pwr.wroc.pl>. Information on using Linux with Polish-specific features. Updated 8 February 1999.

• Portuguese-HOWTO, by Carlos Augusto Moreira dos Santos <casantos@cpmet.ufpel.tche.br>. Este documento pretende ser um guia de referência de configuraçao do Linux e seus programas.... Updated 24 May 1999.

• PostgreSQL HOWTO, by Al Dev (Alavoor Vasudevan) <aldev@hotmail.com>. How to setup PostgreSQL as a database server. Updated 23 October 1999.

• Printing HOWTO, by Grant Taylor <gtaylor+pht@picante.com>. HOWTO on printing software for Linux. Updated 9 March 1999.

• Printing Usage HOWTO, by Mark Komarinski <markk@auratek.com>. How to use the printing system for a variety of file types and options. Updated 6 February 1998.

• Quake HOWTO, by Bob Zimbinski <bobz@mr.net> Thomas Mike Hallock medina.net>. This document explains how to install, run and troubleshoot Quake, QuakeWorld and Quake II on an Intel Linux system. Updated 30 August 1998.

• RPM HOWTO, by Donnie Barnes <djb@redhat.com>. How to use the Red Hat Package Manager (.rpm). Updated 8 April 1997.

• Reading List HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. Interesting books pertaining to Linux subjects. Updated 20 April 1999.

• Root RAID HOWTO, by Michael A. Robinton <michael@bzs.org>. How to create a root-mounted RAID filesystem. Updated 25 March 1998.

• SCSI Programming HOWTO, by Heiko Eissfeldt <heiko@colossus.escape.de>. Information on programming the generic Linux SCSI interface. Updated 7 May 1996.

• SMB HOWTO, by David Wood <dwood@plugged.net.au>. How to use the Session Message Block (SMB) protocol with Linux. Updated 25 June 1999.

• SMP HOWTO, by David Mentré <David.Mentre@irisa.fr>. This HOWTO reviews main issues (and I hope solutions) related to SMP configuration under Linux. Updated 21 October 1999.

• SRM HOWTO, by David Mosberger <davidm@azstarnet.com>. How to boot Linux/Alpha using the SRM firmware. Updated 17 August 1996.

• Security HOWTO, by Kevin Fenzi <kevin@scrye.com>. General overview of security issues. Updated 25 April 1999.

• Serbian HOWTO, by Zoltan Csala <zolika@sezampro.PRVO-SKLONI-OVO.yu>. Information on using Linux with Serbian-specific features. Updated November 1998.

• Serial HOWTO, by David Lawyer <bf347@lafn.org>. This HOWTO covers basic info on the Serial Port and multiport serial cards. Updated October 1999.

• Serial Programming HOWTO, by Peter H. Baumann <Peter.Baumann@dlr.de>. How to use serial ports in programs. Updated 22 January 1998.

• Shadow Password HOWTO, by Michael H. Jackson <mhjack@tscnet.com>. How to obtain, install, and configure shadow passwords. Updated 3 April 1996.

• Slovenian HOWTO, by Primoz Peterlin <primoz.peterlin@biofiz.mf.uni-lj.si>. Information on using Linux with Slovenian-specific features. Updated 15 February 1999.

• Software Building-HOWTO, by Mendel Leo Cooper <thegrendel@theriver.com>. How to build software packages. Updated 27 July 1999.

• Software Release Practice HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. Describes good release practices for Linux open-source projects. Updated 18 September 1999.

• Sound HOWTO, by Jeff Tranter <jeff_tranter@pobox.com>. Sound hardware and software for the Linux operating system. Updated 24 March 1999.

• Sound Playing HOWTO, by Yoo C. Chung <wacko@laplace.snu.ac.kr>. How to play various sound formats under Linux. Updated 24 March 1999.

• Spanish HOWTO, by Gonzalo Garcia Agullo <Gonzalo.Garcia-Agullo@jrc.es>. Information on using Linux with Spanish-specific features. Updated 20 August 1996.

• TclTk HOWTO, by Luca Rossetti <lukaros@tin.it>. This document describes the Linux approach to Tcl, a scripting language. Updated 07 November 1998.

• teTeX HOWTO, by Robert Kiesling <kiesling@terracom.net>. How to install the teTeX package (TeX and LaTeX) under Linux. Updated 9 November 1998.

• Text-Terminal HOWTO, by David S. Lawyer <bf347@lafn.org>. This document explains what text terminals are, how they work, how to install and configure them. Updated August 1999.

• Thai HOWTO, by Poonlap Veeratanabutr <poon-v@fedu.uec.ac.jp>. How to configure Linux for use with the Thai characterset. Updated 4 August 1998.

• Tips HOWTO, by Paul Anderson <paul@geeky1.ebtech.net>. HOWTO on miscellaneous tips and tricks for Linux. Updated June 1998.

• Turkish HOWTO, by Görkem Çetin <gorkem@gelecek.com.tr>. Information on using Linux with Turkish-specific features. Updated 21 September 1999.

• UMSDOS HOWTO, by Jacques Gelinas <jacques@solucorp.qc.ca>. How to install and use the UMS-DOS filesystem. Updated 13 November 1995.

• UPS HOWTO, by Harvey J. Stein <abel@netvision.net.il>. Information on using a UPS power supply with Linux. Updated 18 November 1997.

• UUCP HOWTO, by Guylhem Aznar <guylhem@danmark.linux.eu.org>. Information on UUCP software for Linux. Updated 6 February 1998.

• Unix and Internet Fundamentals HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. Describes the working basics of PC-class computers, Unix-like operating systems, and the Internet in nontechnical language. Updated 25 September 1999.

• User Group HOWTO, by Kendall Grant Clark <kclark@ntlug.org>. Tips on founding, maintaining, and growing a Linux User Group. Updated 24 April 1998.

• VAR HOWTO, by Mr. Poet <poet@linuxports.com>. Listing of Linux value added resellers. Updated 22 August 1999.

• VME HOWTO, by John Huggins and Michael Wyrick <vmelinux@va.net>. How to run Linux on your VMEbus Pentium and other PCI local bus based VMEbus processor designs. Updated 30 July 1998.

• VMS to Linux HOWTO, by Guido Gonzato <guido@ibogfs.cineca.it>. How to move from VMS to Linux. Updated 17 September 1999.

• VPN-Masquerade HOWTO, by John D. Hardin <jhardin@wolfenet.com>. This document describes how to configure a Linux firewall to. Updated 9 September 1999.

• Vim Editor HOWTO, by Alavoor Vasudevan <alavoor@yahoo.com>. This document is a guide to very quickly setup Vim editor on Linux or Unix systems. Updated 14 October 1999.

• Virtual Services HOWTO, by Brian Ackerman <brian@nycrc.net>. How to set up virtual hosting services. Updated 15 August 1998.

• WWW HOWTO, by Wayne Leister <n3mtr@qis.net>. How to set up WWW clients and servers. Updated 21 August 1999.

• WWW mSQL HOWTO, by Oliver Corff <corff@zedat.fu-berlin.de>. How to set up a web server database with mSQL. Updated 17 September 1997.

• XFree86 HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. How to obtain, install, and configure XFree86 3.2 (X11R6). Updated 12 May 1999.

• XFree86 Video Timings HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. How to compose a mode line for XFree86. Updated 13 June 1999.

• X Window User HOWTO, by Ray Brigleb <ray@croftj.net>. Information on configuring the X Window environment for the Linux user. Updated 1 September 1999.

3.2 mini-HOWTOs

The following mini-HOWTOs are available:

• 3 Button Mouse mini-HOWTO, by Geoff Short <geoff@kipper.york.ac.uk>. How to configure your mouse to use 3 buttons. Updated 31 May 1998.

• ADSM Backup mini-HOWTO, by Thomas Koenig <Thomas.Koenig@ciw.uni-karlsruhe.de>. How to install and use the ADSM backup program. Updated 15 January 1997.

• Asymmetric Digital Subscriber Loop (ADSL) mini-HOWTO, by David Fannin <dfannin@dnai.com>. Addresses the ordering, installation and configuration. Updated 10 April 1999.

• Advocacy mini-HOWTO, by Paul L. Rogers <Paul.L.Rogers@li.org>. Suggestions on how to advocate the use of Linux. Updated 7 May 1998.

• Alsa Sound mini-HOWTO, by Valentijn Sessink. Describes the installation of the Alsa sound drivers for Linux. Updated 18 May 1999.

• Apache SSL PHP/FI frontpage mini-HOWTO, by Marcus Faure <marcus@faure.de>. Build a multipurpose webserver. Updated July 1998.

• Automount mini-HOWTO, by Don <don@sabotage.org>. This file describes the autofs automounter, how to configure it, and points out some problems to avoid. Updated 17 April 1999.

• Backup with MSDOS mini-HOWTO, by Christopher Neufeld <neufeld@physics.utoronto.ca>. How to backup Linux machines with MSDOS. Updated 5 August 1997.

• Battery Powered mini-HOWTO, by Hanno Mueller <hanno@lava.de>. How to reduce a Linux system’s power consumption. Updated 21 December 1997.

• Boca mini-HOWTO, by David H Dennis <david@freelink.net>. How to install a Boca 16-port serial card (Boca 2016). Updated 1 August 1997.

• BogoMips mini-HOWTO, by Wim C.A. van Dorst <baron@clifton.hobby.nl>. Information about BogoMips. Updated 8 February 1999.

• Bridge mini-HOWTO, by Chris Cole <cole@lynkmedia.com>. How to setup an ethernet bridge. Updated 7 September 1998.

• Bridge+Firewall mini-HOWTO, by Peter Breuer <ptb@it.uc3m.es>. How to setup and ethernet bridge and firewall. Updated 19 December 1997.

• Bzip2 mini-HOWTO, by David Fetter <dfetter@best.com>. How to use the new bzip2 compression program. Updated 22 August 1998.

• Cipe+Masquerading mini-HOWTO, by Anthony Ciaravalo <acj@home.com>. How to setup a Virtual Private Network between your LAN and other LAN’s using cipe through linux masquerading firewall machines. Updated 28 October 1998.

• Clock mini-HOWTO, by Ron Bean <rbean@execpc.com>. How to set and keep your clock on time. Updated July 1999.

• Coffee mini-HOWTO, by Georgatos Photis <gef@ceid.upatras.gr>. Thoughts about making coffee with Linux (humorous). Updated 15 January 1998.

• Cyrus IMAP mini-HOWTO, by Kevin Mitchell <kevin@iserv.net>. How to install the Cyrus IMAP server. Updated 21 January 1998.

• DPT Hardware RAID mini-HOWTO, by Ram Samudrala <me@ram.org>. How to configure hardware RAID. Updated 9 April 1999.

• Diald mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use ’diald’ to dial an ISP. Updated 3 June 1996.

• Ext2fs Undeletion mini-HOWTO, by Aaron Crane <aaronc@pobox.com>. How to retrieve deleted files from an ext2 filesystem. Updated 2 February 1999.

• Fax Server mini-HOWTO, by Erez Strauss <erez@newplaces.com>. How to setup a fax server. Update 8 November 1997.

• Firewall Piercing mini-HOWTO, by François-René Rideau <rideau@ens.fr>. Using ppp over telnet transparently through an Internet firewall. Updated 27 November 1998.

• GIS-GRASS mini-HOWTO, by David A. Hastings <dah@ngdc.noaa.gov>. How to install Geographic Information System (GIS) software. Updated 13 November 1997.

• GTEK BBS-550 mini-HOWTO, by Wajihuddin Ahmed <wahmed@sdnpk.undp.org>. How to setup the GTEK BBS-550 multiport board with Linux. Updated 20 August 1997.

• Hard Disk Upgrade mini-HOWTO, by Yves Bellefeuille <yan@ottawa.com>. How to copy a Linux system from one hard disk to another. Updated 9 June 1999.

• IO Port Programming mini-HOWTO, by Riku Saikkonen <Riku.Saikkonen@hut.fi>. How to use I/O ports in C programs. Updated 28 December 1997.

• IP Alias mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use IP aliasing. Updated 13 January 1997.

• IP Subnetworking mini-HOWTO, by Robert Hart <hartr@interweft.com.au>. Why and how to subnetwork an IP network. Updated 31 March 1997.

• ISP Connectivity mini-HOWTO, by Michael Strates <mstrates@croftj.net>. How to get mail and news over a dialup connection. Updated 6 November 1997.

• Install From ZIP mini-HOWTO, by Kevin Snively <k.snively@seaslug.org>. How to install Linux from a parallel port ZIP drive. Updated 29 April 1998.

• Kerneld mini-HOWTO, by Henrik Storner <storner@osiris.ping.dk>. How to use ’kerneld’ (dynamic module loading). Updated 19 July 1997.

• LBX mini-HOWTO, by Paul D. Smith <psmith@baynetworks.com>. How to use Low-Bandwidth X (LBX). Updated 11 December 1997.

• LILO mini-HOWTO, by Alessandro Rubini <rubini@linux.it>. Examples of typical LILO installations. Updated 16 August 1998.

• Large Disk mini-HOWTO, by Andries Brouwer <aeb@cwi.nl>. How to use disks with > 1024 cylinders. Updated 12 September 1999.

• Leased Line mini-HOWTO, by Rob van der Putten <rob@sput.webster.nl>. How to setup leased line modems. Updated July 1998.

• Linux+DOS+Win95+OS2 mini-HOWTO, by Mike Harlan <r3mdh@raex.com>. How to use Linux and DOS and OS/2 and Win95 together. Updated 11 November 1997.

• Linux+FreeBSD mini-HOWTO, by Niels Kristian Bech Jensen <nkbj@image.dk> How to use Linux and FreeBSD together. Updated 15 January 1999.

• Linux+NT-Loader mini-HOWTO, by Bernd Reichert <reichert@dial.eunet.ch>. How to use Linux and the Windows NT boot loader together. Updated 2 September 1997.

• Linux+Win95 mini-HOWTO, by Jonathan Katz <jkatz@in.net>. How to use Linux and Windows95 together. Updated 26 October 1996.

• Loadlin+Win95 mini-HOWTO, by Chris Fischer <protek@brigadoon.com>. How to use Linux and Windows95 together, using loadlin. Updated 13 March 1999.

• Loopback-Root-FS mini-HOWTO, by Andrew M. Bishop <amb@gedanken.demon.co.uk>. How to use the Linux loopback device to create a Linux native filesystem format installation that can be run from a DOS partition without re-partitioning. Updated 24 September 1999.

• Mac Terminal mini-HOWTO, by Robert Kiesling <kiesling@terracom.net>. How to use an Apple Macintosh as a serial terminal. Updated 9 November 1997.

• Mail Queue mini-HOWTO, by Leif Erlingsson <leif@lege.com>. How to queue remote mail and deliver local mail. Updated 3 September 1997.

• Mail2News mini-HOWTO, by Robert Hart <iweft@ipax.com.au>. How to set up a mail to news gateway. Updated 1 July 1999.

• Man Page mini-HOWTO, by Jens Schweikhardt <schweikh@noc.dfn.de>. How to write man pages. Updated July 1998.

• Modules mini-HOWTO, by Riley H. Williams <rhw@bigfoot.com>. How to setup and configure kernel modules. Updated 14 November 1997.

• Multiboot using LILO mini-HOWTO, by Renzo Zanelli <rzanelli@southeast.net>. Howto multiboot between Windows 95, Windows NT, and Linux. Updated 26 March 1998.

• NCD X Terminal mini-HOWTO, by Ian Hodge <ihodge@nortel.ca>. Describes how to connect an NCD X terminal to a UNIX host. Updated 3 April 1998.

• NFS-Root mini-HOWTO, by Andreas Kostyrka <andreas@ag.or.at>. How to set up diskless Linux machines. Updated 8 August 1997.

• NFS-Root-Client mini-HOWTO, by Ofer Maor <ofer@hadar.co.il>. How to set up diskless Linux machines using NFS. Updated 02 February 1999.

• Netrom-Node mini-HOWTO, by Karl Larsen <k5di@yahoo.com>. How to setup the ax25-utilities package for Amateur Radio such making Netrom Nodes. Updated 19 October 1998.

• Netscape+Proxy mini-HOWTO, by Sarma Seetamraju <sarma@usa.net>. How to set up a proxy server for Netscape. Updated 15 August 1997.

• Netstation mini-HOWTO, by Kris Buytaert <Kris.Buytaert@advalvas.be>. How to hook up a IBM Netstation to your local network using a Linux box as server. Updated 22 February 1998.

• News Leafsite mini-HOWTO, by Florian Kuehnert <sutok@gmx.de>. How to set up a leaf news site. Updated 4 January 1998.

• Offline Mailing mini-HOWTO, by Gunther Voet <freaker@tuc.ml.org>. How to set up email addresses without a dedicated Internet connection. Updated 4 June 1998.

• PLIP mini-HOWTO, by Andrea Controzzi <controzz@cli.di.unipi.it>. How to set up PLIP (Parallel Line Interface Protocol). Updated 12 March 1998.

• Partition mini-HOWTO, by Kristian Koehntopp <kris@koehntopp.de>. How to choose disk partitions. Updated 3 November 1997.

• Partition Rescue mini-HOWTO, by Rolf Klausen <rolfk@romsdal.vgs.no>. How to rescue deleted Linux partitions. Updated 22 October 1997.

• Path mini-HOWTO, by Esa Turtiainen <etu@dna.fi>. How to use the PATH environment variable. Updated 15 November 1997.

• Pre-installation Checklist mini-HOWTO, by S. Parthasarathy <algolog@hd1.vsnl.net.in>. Preinstallation checklist and questionnaire. Updated 29 August 1998.

• Process Accounting mini-HOWTO, by Albert M.C. Tam <bertie@scn.org>. How to set up process accounting. Updated 8 August 1997.

• Proxy ARP Subnet mini-HOWTO, by Bob Edwards <bob@faceng.anu.edu.au>. How to use proxy ARP with subnetting. Updated August 1997.

• Public Web Browser mini-HOWTO, by Donald B. Marti Jr. <dmarti@best.com>. How to set up a guest account to use a WWW browser. Updated 5 January 1998.

• Qmail+MH mini-HOWTO, by Christopher Richardson <rdn@tara.n.eunet.de>. How to install qmail and MH. Updated 5 March 1998.

• Quota mini-HOWTO, by Albert M.C. Tam <bertie@scn.org>. How to set up disk quotas. Updated 8 August 1997.

• RCS mini-HOWTO, by Robert Kiesling <kiesling@terracom.net>. How to use RCS (Revision Control System). Updated 14 August 1997.

• RPM+Slackware mini-HOWTO, by Dave Whitinger <dave@whitinger.net>. How to install the Red Hat Package Manager (RPM) under Slackware. Updated 13 April 1998.

• RedHat CD mini-HOWTO, by Morten Kjeldgaard <mok@imsb.au.dk> Peter von der Ahé <pahe+rhcd@daimi.au.dk>. How to make your own CDs from the Red Hat Linux distribution equivalent to the ones commercially available from Red Hat. Updated 22 September 1998.

• Remote  Boot  mini-HOWTO,  by  Marc  Vuilleumier  Stückelberg

<Marc.VuilleumierStuckelberg@cui.unige.ch>. How to set up a server-based boot selector. Updated February 1999.

• Remote X Apps mini-HOWTO, by Vincent Zweije <zweije@xs4all.nl>. How to run remote X applications. Updated 14 July 1998.

• SLIP-PPP Emulator mini-HOWTO, by Irish <irish@eskimo.com>. How to use SLIP-PPP emulators with Linux. Updated 7 August 1997.

• Saving Space mini-HOWTO, by Guido Gonzato <guido@ibogeo.df.unibo.it>. Squeezing your Linux installation into the least possible space. Updated 7 April 1999.

• Sendmail Address Rewrite mini-HOWTO, by Thomas Roessler <roessler@guug.de>. Set up sendmail’s configuration file for the home user’s dial-up access. Updated 6 May 1998.

• Sendmail+UUCP mini-HOWTO, by Jamal Hadi Salim <jamal@glcom.com>. How to use sendmail and UUCP together. Updated August 1998.

• Secure POP via SSH mini-HOWTO, by Manish Singh <yosh@gimp.org>. How to set up secure POP connections using ssh. Updated 30 September 1998.

• Small Memory mini-HOWTO, by Todd Burgess <tburgess@uoguelph.ca>. How to run Linux on a system with a small amount of memory. Update 29 October 1997.

• Software RAID mini-HOWTO, by Linas Vepstas <linas@fc.net>. How to configure software RAID. Updated 21 November 1998.

• Soundblaster AWE mini-HOWTO, by Marcus Brinkmann <Marcus.Brinkmann@ruhr-uni-bochum.de>. How to install the Soundblaster AWE 32/64. Updated 11 January 1998.

• StarOffice mini-HOWTO, by Matthew Borowski <mkb@poboxes.com>. Information on installing the StarOffice suite. Updated 2 June 1998.

• Term Firewall mini-HOWTO, by Barak Pearlmutter <bap@cs.unm.edu>. How to use ’term’ over a firewall. Updated 15 July 1997.

• TkRat mini-HOWTO, by Dave Whitinger <dave@whitinger.net>. How to install and use the TkRat mail program. Updated 15 September 1999.

• Token Ring mini-HOWTO, by Mike Eckhoff <mike.e@emissary.aus-etc.com>. How to use token ring cards. Updated 7 January 1998.

• Ultra-DMA mini-HOWTO, by Brion Vibber <brion@pobox.com>. How to use Ultra-DMA drives and controllers. Updated 27 May 1999.

• Update mini-HOWTO, by Stein Gjoen <sgjoen@nyx.net>. How to stay updated about Linux development. Updated 3 February 1998.

• Upgrade mini-HOWTO, by Greg Louis <glouis@dynamicro.on.ca>. How to upgrade your Linux distribution. Updated 6 June 1996.

• VAIO mini-HOWTO, by Hideki Saito <hideki@chatlink.com>. This document explains installation of Linux on Sony VAIO computers. Updated 16 September 1998.

• VPN mini-HOWTO, by Árpád Magosányi <mag@bunuel.tii.matav.hu>. How to set up a VPN (Virtual Private Network). Updated 7 August 1997.

• Visual Bell mini-HOWTO, by Alessandro Rubini <rubini@linux.it>. How to disable audible bells, and enable visual bells. Updated 11 November 1997.

• Linux Modem Sharing mini-HOWTO, by Friedemann Baitinger <baiti@toplink.net>. How to setup Linux as a modem sharing server for other client computers. Updated 2 November 1997.

• WordPerfect mini-HOWTO, by Wade Hampton <whampton@staffnet.com>. How to set up WordPerfect for Linux. Updated 13 August 1997.

• X Big Cursor mini-HOWTO, by Joerg Schneider <schneid@ira.uka.de>. How to use enlarged cursors with XWindows. Updated 11 August 1997.

• XFree86-XInside mini-HOWTO, by Marco Melgazzi <marco@techie.com>. How to convert XFree86 to XInside modelines. Updated September 1997.

• xterm Title mini-HOWTO, by Ric Lister <ric@giccs.georgetown.edu>. How to put strings into the titlebar of an xterm. Updated 7 January 1998.

• ZIP Install mini-HOWTO, by John Wiggins <jwiggins@comp.uark.edu>. How to install Linux onto a ZIP drive. Updated 10 January 1999.

• ZIP Drive mini-HOWTO, by Kyle Dansie <dansie@ibm.net>. Provides a quick reference quide on setting up and using the Iomega ZIP drive with Linux. Updated 10 January 1999.

3.3 Special HOWTOs

The High Availability HOWTO, by Harald Milz <hm@seneca.muc.de> is available at

http://metalab.unc.edu/pub/Linux/ALPHA/linux-ha/High-Availability-HOWTO.html.  It is not

included with the HOWTO collection because it relies on figures and cannot be distributed in all supported formats.

The Graphics mini-HOWTO, by Michael J. Hammel <mjhammel@graphics-muse.org> is available at http://www.graphics-muse.org/linux/lgh.html. It is not included with the HOWTO collection because it needs to use a lot of images, which don’t translate to other output formats.

3.4 Unmaintained HOWTOs and mini-HOWTOs

There are a number of unmaintained documents at ftp://metalab.unc.edu/pub/Linux/docs/HOWTO/unmaintained.

These are kept around since old documentation is sometimes better than none. However, you should be aware that you are reading old documentation.

4 Writing and Submitting a HOWTO

If you are interested in writing a HOWTO or mini-HOWTO, please get in touch with me FIRST!!! at

linux-howto@metalab.unc.edu.

Here are a few guidelines that you should follow when writing a HOWTO or mini-HOWTO:

• Try to use meaningful structure and organization, and write clearly. Remember that many of the people reading HOWTOs do not speak English as their first language.

• If you are writing a HOWTO, you must use the SGML-Tools package, available from http://www.sgmltools.org/, to format the HOWTO. This package allows us to produce LaTeX (for DVI and PostScript), plain text, and HTML from a single source document, and was designed specifically for the HOWTOs. This also gives all of the HOWTOs a uniform look. It is very important that you format and review the output of the formatting in PostScript, plain text and HTML.

• If you are writing a mini-HOWTO, you can either use SGML (as described above) or HTML. If you use SGML for your mini-HOWTO, it will be pubished along with the HOWTOs in LDP books.

• Make sure that all of the information is correct. I can’t stress this enough. When in doubt, speculate, but make it clear that you’re only guessing.

• Make sure that you are covering the most recent version of the available software. Also, be sure to include full instructions on where software can be downloaded from (FTP site name, full pathname), and the current version number and release date of the software.

• Include an FAQ section at the end, if appropriate. Many HOWTO documents need an “FAQ” or “Common Problems” section to cover information which can’t be covered in the regular text.

• Use other HOWTOs or mini-HOWTOs as a model! The SGML source to the HOWTOs is available on Linux FTP sites. In addition, have a look at the LDP Style Guide <http://metalab.unc.edu/LDP/

     HOWTO/LDP-Style-Guide.html> for some guidelines.

• Make sure that your name, email address, date, and a version number is near the beginning of the document. You could also include WWW addresses and a snail mail address if you want. The standard header is:

             Title
             Author's name and email address
             Version number and date

for example:

             The Linux HOWTO Index
             by Tim Bynum
             v2.10.29, 31 July 1997

• Lastly, be prepared to receive questions and comments about your writing. There are several hundreds of accesses to the HOWTO collection every day from around the world!

After you have written the HOWTO, mail it to me. If you used SGML-Tools, simply mail me the SGML source; I take care of formatting the documents. I’ll also take care of archiving the HOWTOs on metalab.unc.edu and posting them to the various newsgroups.

It is important that you go through me when submitting a HOWTO, as I maintain the archives and need to keep track of what HOWTOs are being written and who is doing what.

Then, all you have to do is send me periodic updates whenever appropriate.

5 Copyright

Copyright (c) 1995 - 1998 by Tim Bynum.

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

If you have questions, please contact Tim Bynum, the Linux HOWTO coordinator, at

linux-howto@metalab.unc.edu via email.